onsider the following data representing the price of plasma televisions (in dollars).

11621162, 12211221, 12041204, 13561356, 12521252, 12401240, 12541254, 12531253, 11921192, 13991399, 12091209, 13611361, 13191319, 12331233, 11361136, 12751275, 13711371, 11741174, 14031403, 11811181, 14071407

Copy Data

Price of Plasma Televisions (in Dollars)
Class Frequency Class Boundaries Midpoint Relative Frequency Cumulative Frequency
1126–11751126–1175
1176–12251176–1225
1226–12751226–1275
1276–13251276–1325
1326–13751326–1375
1376–14251376–1425

Step 5 of 7 :  

Identify the midpoint of the sixth class.
Answer #1

Solution:

Given that,

Class Frequency Class Boundaries Midpoint Relative Frequency Cumulative Frequency
1126–11751126–1175 3 1125.5 - 1175.5 1150.5 3 / 21 = 0.14 3
1176–12251176–1225 5 1175.5 - 1225.5 1200.5 5 /21 = 0.24 3 + 5 = 8
1226–12751226–1275 6 1225.5 - 1275.5 1250.5 6 /21 = 0.29 8 + 6 = 14
1276–13251276–1325 1 1275.5 - 1325.5 1300.5 1 / 21 = 0.05 14 +1 = 15
1326–13751326–1375 3 1325.5 - 1375.5 1350.5 3 /21 = 0.14 15 + 3 = 18
1376–14251376–1425 3 1375.5 1400.5 3 / 21 = 0.14 18 + 3 = 21